There are 5 green 7 red balls. Two balls are selected one by one without replacement. Find the probability that first is green and second is red.


1
Expert's answer
2021-11-09T12:59:13-0500

Let A = first ball is green, B = second ball is red.

We need to find to the probability that first ball is green and second ball is red:

"P(A\cap B)=P(A)\cdot P(B|A)"

There are 12 balls and 5 of them are green. So, "P(A)=\frac{5}{12}" .

Since 1 green ball is selected, there are 11 balls and 7 of them are red. So, "P(B|A)=\frac{7}{11}" .


"P(A\cap B)=\frac{5}{12}\cdot \frac{7}{11}=\frac{35}{132}" .


Answer: "\frac{35}{132}" .